Solve equations by factoring
Solve the equation.
x^3 + 13x^2 + 40x = 0

Answer by ewatrrr(24785)


Hi,
Solving equation by factoring
x^3 + 13x^2 + 40x = 0
x(x^2 + 13x + 40)=0
x(x+5)(x+8)=0 Note:SUM of the inner product(5x) and the outer product(8x) = 13x
x = 0
(x+5)= 0
x = -5
(x+8)=0
x = -8